Devin left miami and drove at a speed of 45 mph. Timothy left 1 hour later and drove at a speed of 63 mph. How long will it take Timothy to catch up with Deven?
Answer by stanbon(75887)

Devin left miami and drove at a speed of 45 mph. Timothy left 1 hour later and drove at a speed of 63 mph. How long will it take Timothy to catch up with Deven?
========
Devin DATA:
rate = 45 mph ; time = x hrs ; distance = 45x miles
-------------------------------------------
Tim DATA:
rate = 63 mph ; time = x-1 hrs ; distance = 63(x-1) miles
-----
Equation:
distance = distance
45x = 63x-63
18x = 63
x = 3.5 hrs
============================
Cheers,
Stan H.
===============================